Solve the system of following equation by guass elimination method
3x1+2x2+6x3=2
6x1-7x2-11x3=8
-5x1+9x2+3x3=10
1
Expert's answer
2020-12-01T06:11:17-0500

Let us solve the system of following equation by Guass elimination method:


"\begin{cases}3x_1+2x_2+6x_3=2\\
6x_1-7x_2-11x_3=8\\
-5x_1+9x_2+3x_3=10
\end{cases}"


The augmented matrix  of the systems is the following:


"\left[ 
\begin{array}{cccc} 
3 & 2 & 6 & 2\\
6 & -7 & -11 & 8\\
-5 & 9 & 3 & 10
\end{array}
\right]
 {r_2\to r_2-2r_1\choose r_3\to r_3+2r_1}

\left[ 
\begin{array}{cccc} 
3 & 2 & 6 & 2\\
0 & -11 & -23 & 4\\
1 & 13 & 15 & 14
\end{array}
\right]

 {r_1\to r_1-3r_3}"


"\left[ 
\begin{array}{cccc} 
0 & -37 & -39 & -40\\
0 & -11 & -23 & 4\\
1 & 13 & 15 & 14
\end{array}
\right]

{r_1\to r_1-3r_2}

\left[ 
\begin{array}{cccc} 
0 & -4 & 30 & -52\\
0 & -11 & -23 & 4\\
1 & 13 & 15 & 14
\end{array}
\right]

{r_2\to r_2-3r_1}"


"\left[ 
\begin{array}{cccc} 
0 & -4 & 30 & -52\\
0 & 1 & -113 & 160\\
1 & 13 & 15 & 14
\end{array}
\right]

{r_1\to r_1+4r_2}

\left[ 
\begin{array}{cccc} 
0 & 0 & -422 & 588\\
0 & 1 & -113 & 160\\
1 & 13 & 15 & 14
\end{array}
\right]"


Consequently, we have the following system:


"\begin{cases}-422x_3=588\\
x_2-113x_3=160\\
x_1+13x_2+15x_3=14
\end{cases}"


"\begin{cases}x_3=-\frac{294}{211}\\
x_2=160+113(-\frac{294}{211})\\
x_1=14-13x_2-15(-\frac{294}{211})
\end{cases}"


"\begin{cases}x_3=-\frac{294}{211}\\

x_2=\frac{538}{211}\\

x_1=14-13(\frac{538}{211})-15(-\frac{294}{211})
\end{cases}"


"\begin{cases}x_3=-\frac{294}{211}\\

x_2=\frac{538}{211}\\

x_1=\frac{370}{211}
\end{cases}"